Find the quantum ( photon) energy associated with a 650.kHz AM radio signal.

1
Expert's answer
2022-03-07T11:44:06-0500

The energy in one photon is:


"E = h\nu"

where "h = 6.63\times 10^{-34}J\cdot s" is the Planck's constant, and "\nu = 650kHz = 6.5\times 10^5Hz" is the frequency of the light. Thus, obtain:


"E = 6.63\times 10^{-34}J\cdot s\cdot 6.5\times 10^5Hz = 1.02\times 10^{-29}J"

Answer. "1.02\times 10^{-29}J".